of the vertices of a square a(-2,3)&b(8,5) find the equations of the sides of the square
 

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
side of square from ( – 2,3) and (8, 5) is
 
y – 5 = (1/5) ( x – 8)
 
5y – 25 = x – 8
 
x – 5y + 7 = 0
 
Hope it helps